i have a couple questions
what is the following answers?
(19)6y+4=y+14

Answer by tutorcecilia(2152)

6y+4=y+14 [solve for y]
6y-y+4-4=y-y+14-4
6y-y=14-4
5y=10
5y/5=10/5
y=2
.
check by plugging (y=3) back into the original equation:
6y+4=y+14
6(2)+4=(2)+14
12+4=16
16=16 [checks out]